# Imaging-spectroscopy diagnosis of the giant sloshing spiral in the Virgo cluster with the Einstein Probe Follow-up X-ray Telescope

### v2

@@
We performed deep X-ray observations of the Virgo cluster using the Einstein Probe Follow-up X-ray Telescope (EP-FXT) with a total exposure of 295 ks. Taking advantage of the large field of view (FoV) and low particle background of EP-FXT, the image reveals a giant spiral feature connecting the cold fronts in the northwest and southeast, forming a coherent structure consistent with previous XMM-Newton and Suzaku findings. Furthermore, we present two-dimensional maps of temperature, metallicity, and entropy in the Virgo Cluster, covering a FoV of approximately $28.5^{\prime}$. These maps clearly depict a spiral structure characterized by high density, low temperature, high metallicity, and low entropy. These results support the scenario in which the spiral morphology originates from gas sloshing induced by a minor merger. In addition, the temperatures measured with EP-FXT show reasonable consistency with those obtained from XMM-Newton, within uncertainties.
@@ 1 Introduction
Galaxy clusters, composed of hundreds to thousands of galaxies bound by gravity, are the largest building blocks of the cosmic web (Kravtsov & Borgani 2012; Vikhlinin et al. 2014). The X-ray radiation from clusters predominantly arises from the intracluster medium (ICM), a hot, diffuse plasma where a wealth of complex structures can be seen, such as shocks, bubbles, and cold fronts, which trace the past and ongoing violent dynamic processes in the system (Markevitch & Vikhlinin 2007). Cold fronts, distinguished from shock fronts, are sharp interfaces separating brighter, denser, and cooler gases from the other side with a continuous pressure across the interface (Owers et al. 2009; de Plaa et al. 2010; Ettori et al. 2013; Ghizzardi et al. 2010).
@@
Cold fronts or even spectacular spiral structures have been widely identified in nearby clusters (Keshet 2012). It is proposed that the cold front originates from a displacement of cold gas at the cluster core, triggered by gravitational perturbations—such as those induced by minor mergers (Ascasibar & Markevitch 2006; ZuHone et al. 2010; Rossetti 2013; Markevitch et al. 2001; Roediger et al. 2011; Markevitch & Vikhlinin 2007). In relaxed galaxy clusters, minor mergers disturb the central gas, initiating sloshing motions that produce multiple cold fronts at various locations over time (Roediger et al. 2011; Liu et al. 2018). These sloshing motions can evolve into intricate spiral structures, visible in X-ray images as concentric edges (Ascasibar & Markevitch 2006; Sonkamble et al. 2024), such as the case of A795 (Kadam et al. 2024a). Alternative explanations for the sloshing-induced spiral include two-phase gas flows with differing velocities (Keshet 2012) and the rotational motion of an in-falling gas flow around a moving brightest cluster galaxy (BCG) due to the Coriolis force (Inoue 2022).
@@
Despite the detection of a variety of cold fronts, those manifesting themselves as a clear and complete spiral structure are rare, mainly due to difficulties in revealing large-scale and low-contrast structures with current X-ray telescopes. So far, remarkable examples include those seen in A496, A795, A1775, A2029, A2052, A2566, A2657, Centaurus, NGC 7618/UGC 12491, Perseus, and Virgo (Walker et al. 2017; Fabian et al. 2006; Roediger et al. 2012; Ghizzardi et al. 2014; Kadam et al. 2024a; Botteon et al. 2024; Machacek et al. 2023; Blanton et al. 2011; Paterno-Mahler et al. 2013; Sanders et al. 2016b; Botteon et al. 2021; Kadam et al. 2024b), mainly thanks to mosaic observations with XMM-Newton and Chandra. Detection and detailed investigations of large-scale sloshing spirals are essential for advancing our understanding of the physical mechanism of sloshing and dynamical evolution of galaxy clusters.
@@
Because of its proximity, the Virgo cluster serves as an excellent target for in-depth studies on this topic. Within the Virgo cluster, M87 stands out as the largest and brightest elliptical galaxy, with a redshift of 0.0042. It has been extensively studied across multiple wavelengths, providing valuable insights into the interaction between its central supermassive black hole ($3.2\times 10^{9}M_{\odot}$ within 3.5 pc of its center) and the surrounding ICM (Harms et al. 1994; Simionescu et al. 2010; Bahcall & Sarazin 1977). Early X-ray observations with ROSAT, XMM-Newton, and Chandra have provided definitive evidence for active galactic nucleus (AGN) feedback surrounding the central galaxy M87 (Böhringer et al. 2001; Molendi & Gastaldello 2001; Bohringer et al. 1995; Forman et al. 2005; Sanders et al. 2016a). In addition, several distinct structures within M87 have been discovered, such as two X-ray arms, bubbles, and cavities (Sanders et al. 2016a). X-ray observations have revealed multiple sloshing cold fronts in the Virgo Cluster. XMM-Newton identified cold fronts at 33 kpc and 90 kpc, and revealed a spiral structure extending to approximately 200 kpc (Simionescu et al. 2010; Roediger et al. 2011). Suzaku later detected additional outer cold fronts at 233 and 280 kpc and carried out detailed radial analyses of their thermodynamic properties, further confirming that gas sloshing extends well beyond the core of the cluster (Simionescu et al. 2017). Although previous studies have mapped the two-dimensional thermodynamic properties of the ICM in the central region of M87 Gatuzz et al. (2022), the temperature, entropy, and metallicity distributions in the outer spiral arm remain poorly characterized. While eROSITA is capable of covering the extended spiral structure on the outskirts of Virgo (McCall et al. 2024), its relatively short exposure time has limited detailed analysis of the associated physical properties. To more comprehensively investigate the role of gas sloshing and cold fronts in the outskirts of galaxy clusters, it is important to conduct observations with an X-ray telescope that offers a larger FoV and sufficient sensitivity.
@@
Benefiting from its large FoV and low particle background, the Follow-up X-ray Telescope onboard the Einstein Probe (EP-FXT)(Yuan et al. 2022, 2025) is well suited for detecting large-scale, low surface brightness extended structures. In this paper, we present deep EP-FXT observations of the Virgo cluster. Our data support the detection of the giant sloshing spiral and enable spatially resolved analyses of its thermodynamic properties. The paper is organized as follows. We describe the observations in Sect. 2, with imaging results presented in Sect. 3 and spatially-resolved spectral results in Sect. 4. We discuss the physical implications in Sect. 5 and summarize our findings in Sect. 6.
@@ 2 Observations and data reduction
EP-FXT is one of the two payloads onboard the EP mission,which is an X-ray telescope as a result of China-Europe collaboration, launched on January 9, 2024. EP-FXT is capable of conducting detailed observations of targets triggered by the Wide-field X-ray Telescope (EP-WXT), as well as preselected targets designated from the ground. EP-FXT comprises two identical modules, FXT-A and FXT-B, each equipped with Wolter-I type X-ray optics featuring a focal length of 1.6 m and 54 nested gold-coated nickel mirror shells. These modules are paired with PNCCD detectors as focal-plane instruments, enabling precise measurements of photon energy, position, and timing (Cui et al. 2023). The device features a pixel array of 384 $\times$ 384, with an angular resolution of approximately $9.6^{\prime\prime}$ per pixel. EP-FXT covers an energy range of 0.3–10 keV, with a FoV of $1^{\circ}\times 1^{\circ}$, a spatial resolution of $22^{\prime\prime}$ in terms of half-power diameter (HPD), and an effective area of approximately 300 $cm^{2}$ at 1 keV (one unit). In addition, the particle background of EP-FXT is about one-fifth that of eROSITA, and its low background characteristics make it particularly well-suited for observing low-brightness diffuse emissions (Zhang et al. 2022).
@@
As the first-light object and a calibration source for EP-FXT, the Virgo cluster and its central galaxy M87 have been extensively observed. Due to the large FoV of EP-FXT, a single observation can effectively cover the outskirts of M87. All observations were performed in full-frame (FF) mode, using either the thin or medium filter. Three of the FXT-B observations were conducted in a closed mode, rendering the corresponding data unusable. As a result, only FXT-A data were used, resulting in a total effective exposure time of 295 ks. (see an observational log in Table LABEL:label1).
@@
Figure 1: Combined EP-FXT X-ray image of the Virgo cluster in the 0.3-2.5 keV band. The image has been vignetting corrected, background and point-source subtracted, flattened, and adaptively smoothed. The white circle marks the background region. The bar has a length of 10. The arrows point to north and west.
@@
We used the latest version 1.10 of the Follow-up X-ray Telescope Data Analysis Software (FXTDAS) [^1] http://epfxt.ihep.ac.cn/analysis developed at the EP-FXT Science Center for data reduction (Zhao et al. 2025). New event files were generated through the fxtchain pipeline, and the out-of-time events were identified using fxtootest, which were then removed from the data with ftimgcalc [^2] https://heasarc.gsfc.nasa.gov/docs/software/lheasoft/help/heatools.html. When high-energy cosmic rays pass through space telescopes, they release energy onto the detectors, exciting fluorescence lines that are not modulated by the vignetting effect. To correct for this background contribution, fxtbkggen was used to generate particle backgrounds for each observation. We removed the particle background from each image and then used the reproject_image [^3] Derived from the Chandra Interactive Analysis of Observations (CIAO) version 4.16.0, available at https://cxc.cfa.harvard.edu/ciao/ tool to merge all exposure images along with their corresponding particle-background-subtracted images. Subsequently, we used the merged exposure maps to correct for the vignetting effect in the particle background-subtracted X-ray images. Following this, the data from ObsID 13600005125, with the particle background already subtracted, were selected as the reference for the diffuse background. Given that the primary focus of this study is M87, we selected a background region in close proximity to the M87. This approach reduces the potential to underestimate the diffuse background component surrounding M87, thereby facilitating more precise background subtraction and strengthening the overall robustness of the analysis. We combined multiple observations, subtracted the diffuse background, and normalized the image by the maximum exposure time to produce a calibrated X-ray image. We used wavedetect3 with scales = ”4 8 16 32” and the default sigthresh to detect and remove point sources. The voids were filled using dmfilth3 with the POISSON method and a random seed of 0 to ensure reproducibility. Adaptive smoothing was then applied using dmimgadapt with Gaussian kernels, smoothing scales from 0.1 to 10 pixels, and a minimum of 30 counts per kernel. The final X-ray image in the energy range of 0.3-2.5 keV shown in Fig. 1.
@@
Furthermore, to compare with previous results, we reanalyzed a Chandra observation (ObsID 5826) with an exposure of 126 ks, and an XMM-Newton observation (ObsID 0803670601, EPIC-MOS2) with an exposure of 65 ks. The data were reprocessed using the XMM-Newton Science Analysis System (SAS  [^4] https://www.cosmos.esa.int/web/xmm-newton/sas), and time intervals affected by flares were removed through high-energy light curve screening. In the subsequent XMM-Newton analysis, blank-sky background files were employed to improve the accuracy of background estimation and spectral analysis.
@@ 3 Imaging analysis
To visualize faint structures in the image, we applied two different techniques for image enhancement.
Figure 2: Adaptive GGM filtered image in the central Virgo cluster in 0.3-2.5 keV with SNR = 70. Prominent features are labeled. The two dashed line to mark the giant spiral are identical to those in Fig. 4.
@@
Gradient filtering enhances the visibility of both sharp and flat image features, thereby facilitating their detection and analysis (Sanders et al. 2016b, a, 2022). Because of the high noise in X-ray image gradient measurements within low-brightness regions, the adaptive Gaussian Gradient Magnitude (GGM) filtering is employed for dynamic scale selection. This method dynamically adjusts filter parameters based on local gradients, enabling better capture of image textures and edges (Sanders et al. 2022). Additionally, adjusting the signal-to-noise ratio (SNR) helps smooth the image, enhances the SNR, and reveals finer details. Using these techniques, we applied the adaptive GGM filter to the vignetting-corrected and background-subtracted image in the 0.3–2.5 keV energy range, see Fig. 2.
@@
The image clearly shows two X-ray arms, with the southwestern arm exhibiting a clockwise rotation toward the southeast (Forman et al. 2007). This rotational feature may be attributed to the interaction between relativistic jets from the AGN and the ICM (Forman et al. 2005). Furthermore, the image also identifies a cavity, a shock to the northwest, and cold fronts to the southeast and northwest. These features are consistent with the analysis by Sanders et al. (2016a) using the Chandra data. We also identified a region of significantly reduced surface brightness outward from the southwestern X-ray arm, which may indicate the presence of a cavity. The outskirts of M87 exhibit a prominent spiral morphology, extending from the northwest to the southeast, with similar spiral features also visible in the XMM-Newton mosaic image (Roediger et al. 2011; Simionescu et al. 2010). We then generate a residual image to better visualize this large-scale faint feature.
@@
First, we need to estimate the first-order radial profile using the well-known double $\beta$ model and subtract it. As shown in Fig. 3, the red solid line represents the best-fit double $\beta$ model to the EP-FXT data (black points). The blue points represent the XMM-Newton data and the green points represent the Chandra data, for comparison. As shown, apart from the central region affected by the AGN, the EP-FXT data are in agreement with the double $\beta$ model and exhibit consistency with the radial profiles measured by the three telescopes. Furthermore, thanks to EP-FXT’s large FoV, its observational range extends to regions up to $30^{\prime}$ from the outskirts of M87.
@@
Figure 3: Normalized surface brightness profiles measured with EP-FXT, XMM-Newton, and Chandra. The EP-FXT data are fitted with a double $\beta$ model, with individual components displayed in dashed yellow curves.
@@
Figure 4: Residual image obtained by subtracting the image in Fig. 1 with the elliptical $\beta$ model. A Gaussian smooth with a kernel size of 7 pix is applied (1 pixel = $9.6^{\prime\prime}$). Four regions of interest (NW, SE, SW and NE) are marked. The white dashed lines mark the position of radial discontinuity. The blue dashed lines mark the positions of the giant spiral arm.
@@
To characterize the radial profile, we used Sherpa (version 4.16.0 within CIAO) [^5] https://cxc.cfa.harvard.edu/sherpa/ to model the surface brightness. To account for the apparent elongated morphology as seen from the image, we fitted an elliptical $\beta$ model to the data; the best‑fit parameters are listed in Table 2. We then subtracted this model from the original image to obtain the residual image shown in Fig. 4.
@@
The residual image reveals a few patterns, including X-ray arms and discontinuities which indicate cold fronts or shocks. Notably, we identified a ring-like structure outside the X-ray arms, which displays a break in the southern part. This may be attributed to interactions between the AGN jets and the ICM (Forman et al. 2005). Most importantly, the large-scale spiral structure with excessive brightness is evident on the outskirts, spiraling from the cold front in the northwest to the southern cold gas clump, marked with blue dashed lines in Fig. 4 and white dashed lines in Fig. 2. Consistent with the mosaic XMM-Newton results (Simionescu et al. 2010), the EP-FXT observations clearly reveal the spiral features as continuous and well-defined across the entire FoV, likely owing to the instrument’s low particle background and wide spatial coverage.
@@
The northwestern outer edge (a cold front) of the spiral lies approximately $18.3^{\prime}$ from the center. The structure rotates counterclockwise toward the south, with its southern inner edge located approximately $17.5^{\prime}$ from the center.
@@
Figure 5: An example of the EP-FXT spectrum from a contour-binning region in M87. The black and red points represent the spectral data from FXT-A and FXT-B for ObsID 13600005124, with the corresponding background from ObsID 13600005125 subtracted, and the solid line represents the best-fit model. The green and blue points represent the spectra from FXT-A and FXT-B for ObsID 13600005125, respectively. The light cyan and orange represent the particle backgrounds for ObsIDs 13600005124 and 13600005125, respectively.
@@
Figure 6: Temperature, pressure, entropy, and metallicty maps measured with EP-FXT, pixelated with the contour-binning method. The white dashed curves mark the spiral structure as in Fig. 4.
4 Imaging-spectral analysis

### v3

@@
We performed deep X-ray observations of the Virgo cluster using the Einstein Probe Follow-up X-ray Telescope (EP-FXT) with a total exposure of 295 ks. Taking advantage of the large field of view (FoV) and low particle background of EP-FXT, the image reveals a giant spiral feature connecting the cold fronts in the northwest and southeast, forming a coherent structure consistent with previous XMM-Newton and Suzaku findings. Furthermore, we present two-dimensional maps of the temperature, metallicity, and entropy in the Virgo cluster that cover a FoV of approximately $28.5^{\prime}$. These maps clearly depict a spiral structure that is characterized by high density, low temperature, high metallicity, and low entropy. These results support the scenario in which the spiral morphology originates from gas sloshing that is induced by a minor merger. In addition, the temperatures measured with EP-FXT are reasonably consistent with those obtained from XMM-Newton within the uncertainties.
@@ 1 Introduction
Galaxy clusters are composed of hundreds to thousands of galaxies that are bound by gravity. They are the largest building blocks of the cosmic web (Kravtsov & Borgani 2012; Vikhlinin et al. 2014). The X-ray radiation from clusters predominantly arises from the intracluster medium (ICM), which is a hot, diffuse plasma with a wealth of complex structures, such as shocks, bubbles, and cold fronts that trace past and ongoing violent dynamic processes in the system (Markevitch & Vikhlinin 2007). Cold fronts are distinguished from shock fronts. They are sharp interfaces that separate brighter, denser, and cooler gases from the other side with a continuous pressure across the interface (Owers et al. 2009; de Plaa et al. 2010; Ettori et al. 2013; Ghizzardi et al. 2010).
@@
Cold fronts or even spectacular spiral structures have been widely identified in nearby clusters (Keshet 2012). It is proposed that the cold front originates in a displacement of cold gas at the cluster core, triggered by gravitational perturbations, such as those induced by minor mergers (Ascasibar & Markevitch 2006; ZuHone et al. 2010; Rossetti 2013; Markevitch et al. 2001; Roediger et al. 2011; Markevitch & Vikhlinin 2007). In relaxed galaxy clusters, minor mergers disturb the central gas, initiating sloshing motions that produce multiple cold fronts at various locations over time (Roediger et al. 2011; Liu et al. 2018). These sloshing motions can evolve into intricate spiral structures that are visible in X-ray images as concentric edges (Ascasibar & Markevitch 2006; Sonkamble et al. 2024), for instance, A795 (Kadam et al. 2024a). Alternative explanations for the sloshing-induced spiral include two-phase gas flows with differing velocities (Keshet 2012) and the rotational motion of an in-falling gas flow around a moving brightest cluster galaxy (BCG) due to the Coriolis force (Inoue 2022).
@@
Despite the detection of a variety of cold fronts, those that manifest themselves as a clear and complete spiral structure are rare, mainly because current X-ray telescopes cannot easily reveal large-scale and low-contrast structures. So far, remarkable examples include those seen in A496, A795, A1775, A2029, A2052, A2566, A2657, Centaurus, NGC 7618/UGC 12491, Perseus, and Virgo (Walker et al. 2017; Fabian et al. 2006; Roediger et al. 2012; Ghizzardi et al. 2014; Kadam et al. 2024a; Botteon et al. 2024; Machacek et al. 2023; Blanton et al. 2011; Paterno-Mahler et al. 2013; Sanders et al. 2016b; Botteon et al. 2021; Kadam et al. 2024b), which were mainly found in mosaic observations with XMM-Newton [^1] The X-ray Multi-Mirror Mission and Chandra [^2] Chandra X-ray Observatory. The detection and detailed investigations of large-scale sloshing spirals are essential for advancing our understanding of the physical mechanism of sloshing and of the dynamical evolution of galaxy clusters.
@@
Because of its proximity, the Virgo cluster is an excellent target for in-depth studies on this topic. Within the Virgo cluster, M87 stands out as the largest and brightest elliptical galaxy, with a redshift of 0.0042. It has been extensively studied at multiple wavelengths, which provided valuable insights into the interaction between its central supermassive black hole ($3.2\times 10^{9}M_{\odot}$ within 3.5 pc of its center) and the surrounding ICM (Harms et al. 1994; Simionescu et al. 2010; Bahcall & Sarazin 1977). Early X-ray observations with ROSAT [^3] ROentgen SATellite, XMM-Newton, and Chandra have provided definitive evidence that the feedback of an active galactic nucleus (AGN) surrounds the central galaxy M87 (Böhringer et al. 2001; Molendi & Gastaldello 2001; Bohringer et al. 1995; Forman et al. 2005; Sanders et al. 2016a). In addition, several distinct structures within M87 have been discovered, such as two X-ray arms, bubbles, and cavities (Sanders et al. 2016a). X-ray observations have revealed multiple sloshing cold fronts in the Virgo cluster. XMM-Newton identified cold fronts at 33 kpc and 90 kpc and revealed a spiral structure that extends to approximately 200 kpc (Simionescu et al. 2010; Roediger et al. 2011). Suzaku444originally named Astro-E2 later detected additional outer cold fronts at 233 and 280 kpc and carried out detailed radial analyses of their thermodynamic properties, which further confirmed that gas sloshing extends well beyond the core of the cluster (Simionescu et al. 2017). Although previous studies have mapped the two-dimensional thermodynamic properties of the ICM in the central region of M87 Gatuzz et al. (2022), the temperature, entropy, and metallicity distributions in the outer spiral arm remain poorly characterized. While eROSITA [^5] The extended Roentgen Survey and Imaging Telescope Array is capable of covering the extended spiral structure in the outskirts of Virgo (McCall et al. 2024), its relatively short exposure time has limited a detailed analysis of the associated physical properties. To more comprehensively investigate the role of gas sloshing and cold fronts in the outskirts of galaxy clusters, it is important to conduct observations with an X-ray telescope that offers a larger FoV and sufficient sensitivity.
@@
Benefiting from its large FoV and low particle background, the Follow-up X-ray Telescope on board the Einstein Probe (EP-FXT)(Yuan et al. 2022, 2025) is well suited for detecting large-scale extended structures with a low surface brightness. We present deep EP-FXT observations of the Virgo cluster. Our data support the detection of the giant sloshing spiral and enable spatially resolved analyses of its thermodynamic properties. The paper is organized as follows. We describe the observations in Sect. 2, and we present the imaging results in Sect. 3 and the spatially resolved spectral results in Sect. 4. We discuss the physical implications in Sect. 5 and summarize our findings in Sect. 6.
@@ 2 Observations and data reduction
EP-FXT is one of the two payloads on board the EP mission, which is an X-ray telescope that is a result of China-Europe collaboration and was launched on January 9, 2024. EP-FXT is capable of conducting detailed observations of targets triggered by the Wide-field X-ray Telescope (EP-WXT) and of preselected targets designated from the ground. EP-FXT comprises two identical modules, FXT-A and FXT-B, that are each equipped with Wolter-I type X-ray optics featuring a focal length of 1.6 m and 54 nested gold-coated nickel mirror shells. These modules are paired with PNCCD detectors as focal-plane instruments and enable precise measurements of the photon energy, position, and timing (Cui et al. 2023). The device features a pixel array of 384 $\times$ 384, with an angular resolution of approximately $9.6^{\prime\prime}$ per pixel. EP-FXT covers an energy range of 0.3–10 keV, with a FoV of $1^{\circ}\times 1^{\circ}$, a spatial resolution of $22^{\prime\prime}$ in terms of half-power diameter (HPD), and an effective area of approximately 300 $cm^{2}$ at 1 keV (one unit). In addition, the particle background of EP-FXT is about one-fifth of that of eROSITA, and its low background characteristics make it particularly well suited for observing low-brightness diffuse emissions (Zhang et al. 2022).
@@
As the first-light object and a calibration source for EP-FXT, the Virgo cluster and its central galaxy M87 have been extensively observed. The large FoV of EP-FXT means that a single observation can effectively cover the outskirts of M87. All observations were performed in full-frame (FF) mode using either the thin or medium filter. Three of the FXT-B observations were conducted in closed mode, which rendered the corresponding data unusable. As a result, only FXT-A data were used. This resulted in a total effective exposure time of 295 ks (see the observational log in Table LABEL:label1).
@@
Figure 1: Combined EP-FXT X-ray image of the Virgo cluster in the 0.3-2.5 keV band. The image has been vignetting corrected, background and point-source subtracted, flattened, and adaptively smoothed. The white circle marks the background region. The bar has a length of 10. The arrows point north and west.
@@
We used the latest version 1.10 of the Follow-up X-ray Telescope Data Analysis Software (FXTDAS) [^6] http://epfxt.ihep.ac.cn/analysis developed at the EP-FXT Science Center for data reduction (Zhao et al. 2025). New event files were generated through the fxtchain pipeline, and the out-of-time events were identified using fxtootest, which were then removed from the data with ftimgcalc [^7] https://heasarc.gsfc.nasa.gov/docs/software/lheasoft/help/heatools.html. When high-energy cosmic rays pass through space telescopes, they release energy onto the detectors, which excites fluorescence lines that are not modulated by the vignetting effect. To correct for this background contribution, fxtbkggen was used to generate particle backgrounds for each observation. We removed the particle background from each image and then used the reproject_image [^8] Derived from the Chandra Interactive Analysis of Observations (CIAO) version 4.16.0, available at https://cxc.cfa.harvard.edu/ciao/ tool to merge all exposure images along with their corresponding particle background-subtracted images. Subsequently, we used the merged exposure maps to correct for the vignetting effect in the particle background-subtracted X-ray images. Following this, the data from ObsID 13600005125, without particle background, were selected as the reference for the diffuse background. Because the primary focus of this study is M87, we selected a background region in close proximity to M87. This approach means that the diffuse background component surrounding M87 is probably not underestimated, thereby facilitating preciser background subtraction and strengthening the overall robustness of the analysis. We combined multiple observations, subtracted the diffuse background, and normalized the image by the maximum exposure time to produce a calibrated X-ray image. We used wavedetect8 with scales = ”4 8 16 32” and the default sigthresh to detect and remove point sources. The voids were filled using dmfilth8 with the POISSON method and a random seed of zero to ensure reproducibility. Adaptive smoothing was then applied using dmimgadapt with Gaussian kernels, smoothing scales from 0.1 to 10 pixels, and a minimum of 30 counts per kernel. The final X-ray image in the energy range of 0.3-2.5 keV is shown in Fig. 1.
@@
Furthermore, to compare with previous results, we reanalyzed a Chandra observation (ObsID 5826) with an exposure of 126 ks and an XMM-Newton observation (ObsID 0803670601, EPIC-MOS2) with an exposure of 65 ks. The data were reprocessed using the XMM-Newton Science Analysis System (SAS  [^9] https://www.cosmos.esa.int/web/xmm-newton/sas), and time intervals affected by flares were removed through high-energy light-curve screening. In the subsequent XMM-Newton analysis, blank-sky background files were employed to improve the accuracy of the background estimation and spectral analysis.
@@ 3 Imaging analysis
To better reveal structures in the image, we applied multiple image enhancement techniques to enhance the visibility of both sharp and diffuse edges, thereby highlighting structural details more effectively across different spatial scales.
@@
Gradient filtering enhances the visibility of both sharp and flat image features, thereby facilitating their detection and analysis (Sanders et al. 2016b, a, 2022). Because of the high noise in X-ray image gradient measurements within low-brightness regions, the adaptive Gaussian gradient magnitude (GGM) filtering was employed to select the dynamic scale. This method dynamically adjusts filter parameters based on local gradients, enabling better capture of image textures and edges (Sanders et al. 2022). Additionally, adjusting the signal-to-noise ratio (S/N) helps smooth the image, enhances the S/N, and reveals finer details. Using these techniques, we applied the adaptive GGM filter to the vignetting-corrected and background-subtracted image in the 0.3–2.5 keV energy range (see Fig. 2).
Figure 2: Adaptive GGM filtered image in the central Virgo cluster in 0.3-2.5 keV with S/N = 70. Prominent features are labeled. The two dashed lines that mark the giant spiral are identical to those in Fig. 4.
@@
The image clearly shows two X-ray arms. The southwestern arm exhibits a clockwise rotation toward the southeast (Forman et al. 2007). This rotational feature may be attributed to the interaction between relativistic jets from the AGN and the ICM (Forman et al. 2005). Furthermore, the image also identifies a cavity, a shock to the northwest, and cold fronts to the southeast and northwest. These features are consistent with the analysis by Sanders et al. (2016a) based on Chandra data. We also identified a region in which the surface brightness was significantly reduced outward of the southwestern X-ray arm, which may indicate a cavity. The outskirts of M87 exhibit a prominent spiral morphology that extends from the northwest to the southeast, and similar spiral features are also visible in the XMM-Newton mosaic image (Roediger et al. 2011; Simionescu et al. 2010). We then generated a residual image to better visualize this large-scale faint feature.
@@
First, we estimated the first-order radial profile using the well-known double $\beta$ model and subtracted it (see Fig. 3). Except for the central region that is affected by the AGN, the EP-FXT data agree with the double $\beta$ model and are consistent with the radial profiles measured by the three telescopes. Furthermore, the EP-FXT’s large FoV means that its observational range extends to regions up to $30^{\prime}$ from the outskirts of M87.
@@
Figure 3: Normalized surface brightness profiles measured with EP-FXT, XMM-Newton, and Chandra. The EP-FXT data were fit with a double $\beta$ model. Individual components are displayed as dashed yellow curves.
@@
Figure 4: Residual image obtained by subtracting the image in Fig. 1 with the elliptical $\beta$ model. A Gaussian smoothing with a kernel size of 7 pix was applied (1 pixel = $9.6^{\prime\prime}$). Four regions of interest (NW, SE, SW and NE) are marked. The dashed white lines mark the position of the radial discontinuity. The dashed blue lines mark the position of the giant spiral arm.
@@
To characterize the radial profile, we used Sherpa (version 4.16.0 within CIAO) [^10] https://cxc.cfa.harvard.edu/sherpa/ to model the surface brightness. To account for the apparently elongated morphology in the image, we fit an elliptical $\beta$ model to the data; the best‑fit parameters are listed in Table 2. We then subtracted this model from the original image to obtain the residual image shown in Fig. 4.
@@
The residual image revealed a few patterns, including X-ray arms and discontinuities, which indicate cold fronts or shocks. Notably, we identified a ring-like structure outside the X-ray arms that displays a break in the southern part. This may be attributed to interactions between the AGN jets and the ICM (Forman et al. 2005). Most importantly, a large-scale spiral structure with an excessive brightness is evident in the outskirts, where it spirals from the cold front in the northwest to the southern cold gas clump (marked with dashed blue lines in Fig. 4 and dashed white lines in Fig. 2). Consistent with the mosaic XMM-Newton results (Simionescu et al. 2010), the EP-FXT observations clearly reveal that the spiral features are continuous and well defined across the entire FoV, likely owing to the low particle background and wide spatial coverage of the instrument.
@@
The outer northwestern edge (a cold front) of the spiral lies approximately $18.3^{\prime}$ away from the center. The structure rotates counterclockwise toward the south, and its inner southern edge is located approximately $17.5^{\prime}$ away from the center.
@@
Figure 5: EP-FXT spectrum from a contour-binning region in M87. The black and red points represent the spectral data from FXT-A and FXT-B for ObsID 13600005124, with the corresponding background from ObsID 13600005125 subtracted, and the solid line represents the best-fit model. The green and blue points represent the spectra from FXT-A and FXT-B for ObsID 13600005125, respectively. The light cyan and orange points represent the particle backgrounds for ObsIDs 13600005124 and 13600005125, respectively.
@@
Figure 6: Temperature, pressure, entropy, and metallicty maps measured with EP-FXT, pixeled with the contour-binning method. The dashed white curves mark the spiral structure as in Fig. 4.
4 Imaging spectral analysis